IFFMaster

An IFF Browser Version 1.4

August 11, 1995

Kay Drangmeister

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994, 1995 Kay Drangmeister This document was created using TEX3.14, ‘texinfo.tex’, and ‘makeinfo’.

1 Introduction

IFFMaster is a program that allows you to have a view inside the structure of IFF files. It is not yet another picture viewer or sound sample player, but it displays the internal entities (chunks) of a file. For example, the headers of pictures (ILBM) or sound samples (8SVX) are displayed in clear, so you can directly read the size and depth of the image or the sampling rate.

From time to time you will find an unknown file on your hard disk, let’s say a preference file. Since preference files are often stored in IFFormat, there is a chance you can directy examine the contens. There are some programs which write incorrect IFF files (e.g. the ILBM.CAMG chunk is a likely candidate), or which write additional information (e.g. copyright or author chunks). With IFFMaster you can easily verify such cases.

During the last time I implemented some ways to manipulate files, e.g. delete and move chunks. Chunk contents can be edited using a text or binary editor depending on the content type. In future versions (see Appendix A [History], page 11) there will be options to add new chunks like a DPI chunk for images.

Current features include:

• MUI application

• localized GUI (with AmigaOS 2.1 or later)

• Extensive chunk type library (currently 46 form types, 257 known chunks, 79 of these with comprehensive structure description)

• Chunk contents are presented alternatively as structure, text or hex dump

• Bit fields and enumeration types are displayed in clear

• Fixed point values are printed in decimal (e.g. 8SVX.VHDR.Volume)

• Callback hooks for special attributes, e.g. the Mode-ID inside the CAMG chunk is de-referenced

    (e.g. ‘PAL: Hires’).

IFFMaster is distributed under the concept of freeware. Standard disclaimer applies to this program.

2 Installation

Nothing to it: just leave all files and the catalogs directory as they are in one directory, or alternatively copy the appropriate ‘iffmaster.catalog’ to into the system’s locale directory (i.e. ‘LOCALE:Catalogs/yourlanguage/’). By the way: if you mixed up catalog files for different languages, just use IFFMaster to discover the catalog’s language…:)

System requirements are:

• AmigaOS 2.0 (V37)

• AmigaOS 2.1 (V38) for localized GUI

• AmigaOS 3.0 (V39) for some extended features, like displaying colors in ILBM/CMAP chunks.

• MUI version 2.1 (‘muimaster.library’ v8), See Section 5.1 [MUI], page 10.

If you wish to start IFFMaster inside the user-startup, you should do this with ‘runback IFFMaster ICONIFIED’. This will cause IFFMaster to start as an AppIcon, so that Icons may be dragged onto this Icon.

3 Usage

3.1 Main Window

The main window consists of 3 groups. The topmost one is the file specification group, the middle part ist the chunk list and below there are the action buttons.

The current version has some capabilities to manipulate files, but adding chunks is still not implemented. So the ‘Add’ button is always inactive. However, moving (‘Up’, ‘Down’), deleting (‘Del’), and editing (‘Edit’) chunks is possible. Be warned that files may become unreadable to some applications if you delete mandatory chunks, i.e. if you delete the ILBM/BMHD chunk of a picture it will become undecodeable. To activate the manipulation buttons select ‘Prefs/Editable file’ from menu. If that menu item is inactive (it can be locked if you are still unsure about the program’s functions) you can activate it by opening the preferences window and setting the ‘File

editable switch’ to be ‘off for new files’ or to be ‘left unchanged’, see Section 3.3.2 [Safety],

page 6.

3.1.1 File Specification

There are different ways to open a file. The most common one is by selecting ‘Project/Open...’ from the menu, or by clicking on the popup gadget beside the string gadget for getting a file requester. Alternatively, you can simply drag an icon on IFFMaster’s main window (AppWindow). Furthermore, you can open the file which is currently in the clipboard by using the menu item

‘Project/Open Clip’.

3.1.2 Chunk List

The chunk list can show several types of information: the left column shows the chunk ID (e.g. BMHD), then there is the chunk type (e.g. ILBM), and at the right side there is the size of the chunk. Below the action buttons there is a cycle gadget labeled ‘Show’ that specifies the format of the chunk list. If it states ‘Description’ you will get a description of the chunk’s purpose (e.g. ‘Bitmap Header’), and ‘Contents’ will display a short summary of the chunk’s contents. You can advance the cycle gadget by pressing the SPACE key.

Because it is impossible to display the entire contents of a chunk in a single line you can get a comprehensive list by first selecting a chunk in the chunk list and then pressing ‘Info’ (you can as well double-click on the chunk). A new window containing the chunk’s contents will appear. You can display the contents either as structure, plain text or as a hex dump. Change the presentation by using the register above the contents list, or by pressing CURSOR LEFT and CURSOR RIGHT.

You need not close the contents window to display the contents of another chunk. Just click on that chunk in the chunk list or press CURSOR UP or CURSOR DOWN after having activated the chunk list via the TAB key. The new contents are displayed using the appropriate data type.

3.1.3 Action Buttons

Each of the buttons in this group performs an action on the currently active chunk, i.e. the on that is marked by the cursor in the chunk list.

The buttons that are used to modify a file are inactive if the menu item (switch) ‘File Editable’ is turned off. This is for safety reasons, so that the file structure or contents cannot be damaged unitentionally. Depending on the experience of the user there are different locking modes for that switch, see Section 3.3.2 [Safety], page 6.

Info The contents window will be opened, and additional information about the chunk is desplayed in it.

Edit Depending on the chunk’s contents (text or binary data) a text or binary editor is called to let you edit the chunk’s contents. See Section 3.3.1 [General], page 6.

Add This action is not implemented yet due to extremely low user feedback. So this button is always disabled.

Delete The active chunk is entirely deleted.

Up The chunk is moved before its predecessor. Because chunks can not be moved out of their containers, this action will only succed if the preceeding chunk is not one of the type FORM, CAT or LIST. Furthermore, only chunks containing data can be moved, i.e. container chunks (type FORM, CAT or LIST) are not moveable.

Down The chunk is moved behind its successor. The same restrictions as in ‘Up’ apply.

Top The chunk is moved up as far as possible. The same restrictions as in ‘Up’ apply.

Bottom The chunk is moved up as down as possible. The same restrictions as in ‘Up’ apply.

3.2 Contents Window

The contents window displays the contents of the active chunk, see Section 3.1.2 [Chunk List], page 3. Depending on the type of the contents one of the following three display types is (automatically) used.

Structure This is the most sophisticated method to display the chunk’s contents. All parts of the content are decoded and displayed line by line together with a description. Example: A sampled sound (FORM 8SVX) has a header chunk (VHDR) that is 20 bytes long. Displayed as a sequence of hex numbers it would read: 00021432 00000000 00000020 41560100 00010000. Displayed as a structure it reads like this:

                                 OneShot HiSamples  136,242
                                  Repeat HiSamples  0
                                 Samples / HiCycle  32
                                       Samples / s  16,726
                                         # Octaves  1
                             Compression Technique  None
                                      Volume [0,1]  1.00000

Text Some chunks contain plain text, e.g. ANNO chunks. Texts are formatted blockjustified, but no other processing is done.

Hex Dump If a chunk contains neither structured data nor text the contents are displayed using a hex dump. There are some ways to customize the appearence of the dump, see Section 3.3.4 [Hexdump], page 7.

You can also switch between some display styles by hand, using the register gadgets. For example you can have the above mentioned VHDR chunk displayed as a hex dump instead of a structure. To switch between the display styles you can also use the CURSOR LEFT and CURSOR RIGHT keys, even if the active window is the main window and not the contents window. This way you can control all display functions from the main window.

3.3 Preferences Window

The preferences are grouped on four pages: General, Safety, Layout and Hexdump. They are described in the following sections.

After having edited the preferences you may choose to Save For storing the settings on disk. All future invocations of IFFMaster will use

that settings.

Use For storing the settings in RAM only, so they will only last until you reboot your computer.

Cancel For aborting the adjustments you made and use the previously stored settings.

Closing the window with the close gadget has the same effect as klicking on ‘Cancel’.

3.3.1 General

On this page you can specify the editors.

Text Editor specifies the editor that is used for chunks containing plain text, e.g. ANNO-Chunks. You have to make sure that the editor does not spawn itself from the shell, but runs synchronously. If you are using e.g. the CygnusED you may enter ‘ed -sticky’. If you leave this gadget empty the editor specified in the environment variable EDITOR is used.

Binary Editor specifies the editor that is used for chunks containing binary data. A common name for these editors is file zapper, and are likely to be available on the aminet. This editor has also to run synchronously. If you leave this gadget empty the editor specified in the environment variable BINEDITOR is used.

3.3.2 Safety

The following settings are used to prevent inintentional damage to files. Since numerous ‘okay to do...?’ requests by the program are evenly unacceptable you can adjust between the two extremes safety and ease of use. Beginners and casual users should always use the safe settings.

File Editable switch

In the menu ‘Settings’ there is a switch ‘File Editable’ which turns on or off the file manipulation gadgets, see Section 3.1.3 [Action Buttons], page 4. You can set here the behaviour of that switch.

off and locked is the safest setting. The switch is off and disabled, effectively preventing you from accidentially turing on the editing buttons.

off for new files

turns off the manipulation gadgets each time a new file is opened. If you want to edit files only infrequently this is a good setting.

left unchanged

leaves the switch always in the state you switched it to. If you want to edit many files in a row you have to turn on the manipulation gadgets only once.

Overwrite files

If you want to edit many files without keeping backups you can turn the prompting off. Otherwise it is recommended to leave this gadget always on the safe ‘prompt’ position, because since IFFMaster can be fully driven by keyboard it just needs two keypresses like DEL and S (for ‘Save’) to possibly destroy a file.

3.3.3 Layout

The layout page currently features only one setting:

Hex indicator lets you choose your favourite pre- or postfix string to indicate hexadeximal numbers in the structure page of the contents window. For an example file which contains hex numbers open ‘ENV:Sys/locale.prefs’ and select the LCLE chunk. The first 4 numbers (16 Bytes) are reserved and displayed as hex numbers. NB: This setting does not affect the hex dump, see Section 3.3.4 [Hexdump], page 7.

3.3.4 Hexdump

This page contains gadgets that affect the appearence of the hex dump in the contents window.

Limit # bytes Building hex dumps of very large chunks (e.g. ILBM.BODY) can take a long time. Therefore the number of bytes in a hex dump can be limited to a certain ammount. The default limit is 512 bytes. Since in most cases hex dumps are not very meaningful you may wish to leave this value reasonably low. To adjust the value you man use the slider or the string gadget to the right of it. With the check mark button you can turn the limit off, but this is not recommended.

Offset format The leftmost column of the hex dump shows the offset of the first byte in each row. Use this gadget to specify if that offset should be displayed in deximal or hexadecimal.

Dump characters

If the (fixed-width) font used in the hex dump contains all 256 characters you can turn on to print even normally non-printable characters like LineFeed, 0x0A, which will then be displayed as an inverse ‘J’ or something like that. If only a rectangle appears then your font supports only printable characters and you will have to turn on ‘printable only’. Then all non-printable characters are displayed as a dot. A good way to test this feature is to load a 24 bit ILBM picture. These pictures normally contain three CLUT chunks. Normally these chunks contain all byte values in increasing order.

4 Problems?

Building hex dumps of very large chunks (e.g. ILBM.BODY) can take a long time. Therefore the number of bytes in a hex dump can be limited to a certain ammount. The default limit is 512 bytes, it can be adjusted or turned off in the Preferences, see Section 3.3.4 [Hexdump], page 7.

Some chunks (e.g. FORM) are container chunks with no contents, and therefore there is no presentation. Some other chunks (e.g. BODY) contain raw data, presentation of these chunks is limited to hex dump.

5 Credits

I like to thank:

Stefan Stuntz for MUI, see Section 5.1 [MUI], page 10. Eric Totel for his great MUIBuilder, and for keeping MUIBuilder up with

the needs of IFFMaster :). This program would not exist without it.

Francesco Dipietromaria for the italian translation Kai Iske and Walter Dörwald

for support and betatesting

Thomas Reinhardt for his IFF files H. Phil Duby, Bryan Ewert and Klaus Seistrup

for bug reports and comments

5.1 MUI

                               This application uses
                             MUI - MagicUserInterface
                     (c) Copyright 1993/94 by Stefan Stuntz
     MUI is a system to generate and maintain graphical user interfaces. With
     the  aid  of  a  preferences program, the user of an application has the
     ability to customize the outfit according to his personal taste.
     MUI is distributed as shareware. To obtain a complete package containing
     lots of examples and more information about registration please look for
     a  file  called  "muiXXusr.lha"  (XX means the latest version number) on
     your local bulletin boards or on public domain disks.
               If you want to register directly, feel free to send
                              DM 30.-  or  US$ 20.-
                                       to
                                  Stefan Stuntz
                             Eduard-Spranger-Straße 7
                                  80935 München
                                     GERMANY

Appendix A History

Version 1.0

First release Version 1.1

FIX version 1.0 was compiled with option 68030 - sorry! NEW many new chunk types NEW menu (moved ‘about’ and ‘prefs’ into the menu, gaining some space) NEW specifying file names at command line NEW loading files directly from clipboard NEW improved keyboard handling

Version 1.2 NEW fully localized GUI NEW chunk contents can be saved NEW AppWindow and AppIcon, command line option iconified NEW improved formatting capabilities for structure entries NEW sophisticated IFF parser, even brain-dead formats like EMOD.EMIC are handeled

appropriately NEW chunks now have nominal length, a message is displayed if a chunk is too short or

too long BUG Beta testers reported problems with de-iconify function when an icon was dropped

onto the AppIcon. I removed the said call, so you have to double-click the AppIcon after dropping an icon.

Version 1.3

FIX prefs window: close gadget now functioning, see Section 3.3 [Preferences Window],

page 5. FIX cycle chain (TAB) now works in all windows

NEW new chunk IDs: IAND, IANM, DR2D, RGB8, RGBN, SPLT

NEW hex numbers get a user-chosen indicator, construction of hex dumps is now more

than 3 times faster, see Section 3.3.4 [Hexdump], page 7. NEW IFFs can be saved, even nested FORMs, e.g. images inside animations NEW implemented some editing capabilities (delete chunks), see Section 3.1.3 [Action

Buttons], page 4.

NEW preferences can be saved, see Section 3.3 [Preferences Window], page 5. Version 1.4

NEW Completely changed to GNU-C. Differences are: new startup code, no more ANSI-C functions like sprintf() used, enabling some things like localized output of floating point numbers.

FIX should finally run without ‘locale.library’

NEW Online help

NEW new chunk IDs: AMBA, AMDE, AMIN, AMHU, AMUN, EQE1, RESO, VARS, ...

NEW complete structure descriptions of system preference files, some new hooks. Be sure

      to try out e.g. ‘IFFMaster ENV:sys/locale.prefs’.

NEW color descriptions (RGB values) are now also displayed as a colorfield, if you

      have WB 3.0. Check out ‘IFFMaster ENV:sys/palette.prefs’ or any ILBM/CMAP.

NOTE: You need some free pens for this feature, so you may need to specify a deep screen for IFFMaster inside MUIPrefs.

FIX ‘<Clipboard>’ is now accepted as a name for the clipboard. So ‘Save’ now works as ‘Save Clip’ when working on a clipboard file. As a side-effect you can now enter ‘IFFMaster "<Clipboard>"’ at the command line to load directly from the clipboard.

FIX button key definitions did overlap

NEW ’non-printable’ characters can now be displayed in the hex dump, if the font comprises 256 characters, see Section 3.3.4 [Hexdump], page 7. (Requested by Walter Dörwald)

NEW some more editing capabilities (move chunks), see Section 3.1.3 [Action Buttons], page 4.

NEW Chunks containing text or binary data can be edited via text or binary editors, see Section 3.1.3 [Action Buttons], page 4.

NEW new icon, which is also used as AppIcon now

I would greatly appreciate new ideas and enhancement requests. If you discover any chunk unknown to IFFMaster, please let me know (i.e. try to send me that (short) file or even a description of the chunk’s structure, if known). If you dare to translate the catalog to your language, please have a look at ‘translators.readme’ first.

Appendix B Chunks

Following chunks are currently known to IFFMaster:

[anywhere]
     CSET, FVER, ANNO, AUTH, CHRS, hlid, NAME, TEXT, (c)
3DDD

obje

8SVX
     ATAK, body, CHAN, FADE, PAN , RLSE, SEQN, VHDR
ANIM

ansq

AVCF

avfh, gdat

AVCO

cdat, flag, imag

AVEV

acts, cdat, flag, imag, pars, refl

CDVR

vars

COPR

copi, wait, move

CTLG

LANG, strs

DECK

reso

DEEP

dbod, dgbl, dloc, dpel

DR2D
     ATTR, CMAP, CPLY, DASH, DRHD, FONS, OPLY
DTYP

dthd, dtcd

EMOD

8smp, EMIC, patt

ENVL

LFOI, coen, voen, pien, noen, reen, phen

EQED

eqe1

FAX3

fxhd, gphd, page

FAXX

fxhd, gphd, page

FTXT

fons

GXGA

gada

GXMN

meda

GXUI

ggui

GXWD

wdda

IAND

bpct, CMAP, body

IANM
BMHD, CAMG, body
ILBM

ANHD, asdg, bhcp, bhsi, BMHD, body, CAMG, CCRT, CLUT, CMAP, CRNG, DEST, dlta, dmmy, DPAN, DPI , dpps, DPPV, drng, eqe1, FFEX, fitr, fxd2, fxpl, GRAB, junk, mand, SPRT

ISTG

maxf, sobj

KCXM

VERS, pref

LWOB

pnts, pols, SRFS, surf

MCXB
PREF, VERS
MCXP

pref, vers

MTRX

arry, body, dtyp, stru

PREF

alrt, AMBA, AMDE, AMIN, AMHU, AMUN, conf, CTRY, dfss, evnt, etxt, flop, FONT, gena, genc, gtco, gui, ICTL, INPT, JFIF, keys, LCLE, menu, midi, oper, OSCN, PALT, path, pdat, PGFX, pntr, PRHD, pspd, ptrn, PTXT, punt, SCRM, SERL, shmn, sond, tmac, tmdo, tmex, tmic, tmim, tmmo, tmso, VERS, wbcf, wbpc, xdos

PTCH

inpf, outf, pseq, vers

REAL

rani, ratt, rmtr, robj, rscr, rset, rvrs, rwin

RGB8
BMHD, body, CAMG, CMAP, imrt
RGBN
BMHD, body, CAMG, CMAP, imrt
SC3D

edge, face, hier, lamp, lnam, obsv, path, vert, vnam, wrld

SMUS
SHDR, INS1, inst, snx1, TRAK
SPLT

INFO, body

TACF

tpar, tpbr, tpca, tpcm, tpma, tpp1, TPPA, tppx, tpsc, tpse, VERS

TAKE

tfrm, thdr

TDDD

info, obj

TERM

clip, COMD, cptr, date, dial, emln, fast, FILE, misc, MODM, PATH, PHON, recv, SCRN, send, SERL, SOUN, spek, trml, trns, VERS, winf, wind, XFER

TVP2
TVRX
VILL

crc , mode, moni, VER

I am still looking for descriptions for the chunks in smaller print.

Appendix C Author

Kay Drangmeister Am Hubland 16, Apt. 1323 97074 Würzburg Germany T. ++49-931-7059961 (night+day) K.Drangmeister@insider.sub.de drangmei@informatik.uni-wuerzburg.de Index 17

Index

8

8SVX.VHDR.Volume . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

A

action buttons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Action Buttons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Add . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 4 altering files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 AppIcon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 AppWindow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Author . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

B

Bottom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

C

Chunk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Chunk contents as hex dump. . . . . . . . . . . . . . . . . . . . . . . . 5 Chunk contents as structure. . . . . . . . . . . . . . . . . . . . . . . . . 5 Chunk contents as text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 chunk ID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Chunk List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 chunk list format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 chunk liste . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 chunk type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Chunks known . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Chunks without contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Container chunks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 contents of chunk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Contents Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Credits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 cursor left . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 cursor right . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

D

Del . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 description. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Down . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 4

E

Edit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 4 EMOD.EMIC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 ENV:Sys/locale.prefs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Environment variable: BINEDITOR . . . . . . . . . . . . . . . . 6 Environment variable: EDITOR . . . . . . . . . . . . . . . . . . . . 6

F

File editable switch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 file manipulation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 file specification. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 File Specification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 file zapper . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 format of chunk list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

G

General Preferences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

H

hex dump. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Hexdump Preferences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

I

iconified . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 ILBM.BODY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ILBM.CAMG.ModeID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 ILBM.CLUT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

L

Layout Preferences. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 list format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

M

Main Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 manipulation of files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Index 18

MUI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

P

PREF.LCLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Preferences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Preferences Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Preferences: General . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Preferences: Hexdump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Preferences: Layout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Preferences: Safety. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Prefs/Editable file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 presentation of chunk contents . . . . . . . . . . . . . . . . . . . . . . 4 Problems? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

S

Safety Preferences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Show . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

T

TAB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Top . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

U

Up. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 4 Usage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Table of Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3.1 Main Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.1.1 File Specification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.1.2 Chunk List. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.1.3 Action Buttons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3.2 Contents Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.3 Preferences Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3.3.1 General . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.3.2 Safety. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.3.3 Layout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.3.4 Hexdump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

4 Problems?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 5 Credits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

5.1 MUI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

Appendix A History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Appendix B Chunks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Appendix C Author. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17